The demand functions for each of two goods depend on the prices of the goods, p1 and p2: Q1 = 15 -3p1 + p2 and Q2 = 6 -2p2 + p1. However, each supply curve depends on only its own price: Q1 = 2 + p1 and Q2 = 1 + p2. Solve for the equilibrium: p1, p2, Q1, and Q2
1
Expert's answer
2021-10-11T17:05:44-0400

Demand: "Q1 = 15 -3p1 + p2"

Demand: "Q2 = 6 -2p2 + p1"

Supply: "Q1 = 2 + p1"

Supply: "Q2 = 1 + p2"

At equilibrium, demand = supply

"15 -3p1 + p2 = 2 + p1\\
13 = 4p1 - p2 --------> equation(1)"


At equilibrium, demand = supply,

"6 -2p2 + p1 = 1 + p2\\
5= -p1+ 3 p2 --------> equation(2)"

Solving the two equation gives:

"p2 = 4p1 - 13\\
5 = -p1 + 3(4p1 -13)\\
5 = -p1 + 12p1 - 39\\
44 = 11p1\\
p1 = 4\\ p2 = 4 \times 4 - 13\\
p2 = 3"


Supply, "Q1 = 2 + p1\\"

"Q1 = 2 + 4
\\Q1 = 6\\
Q2 = 1 + p2\\
Q2 = 1 + 3 \\Q2 = 4"

"p1 = 4,\ p2 = 3,\ Q1= 6,\ Q2 = 4"